OpenStudy (anonymous):

Triangle ABC has the given area, angle and side. Find the length of the requested side. Area = 7.39 A=38* b=4 find c

OpenStudy (raden):

use the formula : area = 1/2 bc sin(A)

OpenStudy (raden):

so, 1/2 * 4 * c sin(38) = 7.39 solve for c

OpenStudy (anonymous):

alrighty thanks man, doing a final review packet and that formula slipped through my mind

OpenStudy (anonymous):

what would be the answer btw so i can double check my work

OpenStudy (raden):

1/2 * 4 * c sin(38) = 7.39 2 c (0.615) = 7.39 1.23 c = 7.39 c = 7.39/1.23 = ....

OpenStudy (anonymous):

Alright thanks a ton man i solved to 6